OpenStudy (anonymous):

what's the answer to this one using the elimination method Please? I am so tired of looking at all these numbers until I am confused now! 2x+2y=42 x=y+7 I really

OpenStudy (psymon):

Well, start off by rearranging the bottom equation so that x's, y's, and numbers all line up in both equations.

OpenStudy (anonymous):

My mind is fired and I don't even know where to start

OpenStudy (psymon):

Thats fine. Do you know how to rearrange the bottom equation so that both equations have x's, y's, and numbers lined up?

OpenStudy (anonymous):

x=y+7 x-7=y

OpenStudy (psymon):

I'll help ya out with lining them up to start then. 2x + 2y = 42 x - y = 7 Do you see how I got that?

OpenStudy (anonymous):

You switched the x and changed the sign

OpenStudy (psymon):

I just subtracted y from both sides really :P So now we need something to be eliminated. In order for this to happen, we have to have the same x term in both equations, but opposite signs. We can also have the same y term in both equations that have opposite signs. So as of now, you have 2x + 2y = 42 x - y = 7 So looking at this, if something is going to cancel, I need to have either a 2x in one equation and -2x in the other. Or I could have a 2y in one equation and a -2y in the other. Do you know how this can be done?